# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
$G$ $:=\widetilde{S}_{n},$
$L$ $:=\widetilde{S}_{r,p^{d}},$
$N^{\rho,d}$ $:=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}).$
Note that $N^{\rho,d}$ must permute $V_{1},\cdots,V_{d}$. In particular, $N^{\rho,d}\leq\widetilde{S}_{n}$ is the subgroup generated by $L=\widetilde{S}_{U,V_{1},\cdots,V_{d}}$ and the set $\{T_{w}\mid w\in S_{n}\}$ where the notation $T_{w}$ is from 5.12. Recall the notation in (5.1) that we will also use the subgroups $G_{\bar{0}}=\widetilde{A}_{n}$, $N_{\bar{0}}^{\rho,d}=N^{\rho,d}\cap\widetilde{A}_{n}$, etc. We make the following identifications using (5.3):
$\begin{split}RLe_{z}&\cong R\widetilde{S}_{U}e_{z}\otimes_{R}R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}.\end{split}$ (5.7)
@@
By Proposition 5.9, we have $f\in R_{p}L\subseteq R_{p}N^{\rho,d}$. By [22, page 128], the isomorphism
$RLe_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z})^{\otimes d}$
@@
$RN^{\rho,d}e_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z}\wr_{s}\mathcal{T}_{d}^{R}),\ T_{i}e_{z}\mapsto\kappa_{i}^{-1}(1\otimes\tau_{i}),$
where $\kappa_{i}$’s are from [22, Lemma 4.7]. More specifically, according to the proof of [22, Lemma 4.7], if $\widetilde{S}_{n}=\widetilde{S}_{n}^{+}$, then $\kappa_{i}\in\{1,-1\}$ for all $1\leq i\leq d-1$, and if $\widetilde{S}_{n}=\widetilde{S}_{n}^{-}$, then $\kappa_{i}\in\{\sqrt{-1},-\sqrt{-1}\}$ for all $1\leq i\leq d-1$. In particular, we have the isomorphisms of superalgebras
$\begin{split}RLf&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1})^{\otimes d}\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}\\ RN^{\rho,d}f&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R})\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R}).\end{split}$ (5.8)
@@
For the group $\widetilde{S}_{n}^{+}$, since $\kappa_{i}=\pm 1\in R_{p}$ and since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
For the group $\widetilde{S}_{n}^{-}$, $\kappa_{i}=\pm\sqrt{-1}$. Then $\kappa_{i}\in R_{p}$ if and only if $p\equiv 1~{\rm mod}~4$. Hence if $p\equiv 1~{\rm mod}~4$, the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
Lemma 5.15 ([22, Lemmas 4.9 and 4.10]).
Let $0<d<p$. As in 5.12, for $i=1,\cdots,d$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $i=1,\cdots,d$ and $w\in S_{d}$, and let
$D:=D_{1}\times\cdots\times D_{d}\leq\widetilde{S}_{V}.$
@@
If $\rho$ is a $d$-Rouquier $\bar{p}$-core, we refer to $B^{\rho,d}_{R}$ as an RoCK block of weight $d$ over $R$.
For more information on RoCK blocks we refer to [22, §5].
@@
As in Section 5, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. In this section we continue to adopt the notation in 5.13; we always assume that $d>0$. Let $D$ be a Sylow $p$-subgroup of $\widetilde{S}_{V}$. Then by Theorem 5.6, $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RGb=R\widetilde{S}_{n}e_{\rho,d}$, and by [21, Theorem 5.2.10], $RN_{\widetilde{A}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RG_{\bar{0}}b=R\widetilde{A}_{n}e_{\rho,d}$. Note that we interpret $e_{\varnothing,0}$ as $e_{z}$. We set
$\begin{split}b^{\rho,d}_{R}&:=RN_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R})_{\bar{0}}:=RN_{\widetilde{A}_{n}}(D)e_{\rho,0},\\ b^{\rho,d}_{R_{p}}&:=R_{p}N_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}N_{\widetilde{A}_{n}}(D)e_{\rho,0}.\end{split}$
Theorem 6.1 ([21, Proposition 5.2.13]).
Suppose that $0<d<p$. Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $k$-superalgebras
$b^{\rho,d}_{k}\cong k\widetilde{S}_{U}e_{\rho,0}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}).$ (6.1)
@@
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Suppose that $0<d<p$. Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$b^{\rho,d}_{\mathbb{F}_{p}}\cong\mathbb{F}_{p}\widetilde{S}_{U}e_{\rho,0}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$ (6.2)
@@
$\begin{split}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}&\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\\ \mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}&\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\end{split}$ (6.3)
We realise [21, Lemma 5.2.19] over $\mathbb{F}_{p}$:
@@
Let $P$ be a Sylow $p$-subgroup of $\widetilde{A}_{p}$. There is a splendid Rickard complex of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules that extends to a complex of $(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}}$-modules.
Proof. In [19, Theorem 1.10], Kessar and Linckelmann construct a splendid Rickard complex for blocks with a cyclic defect group over a non-splitting field, hence a splendid Rickard complex $X$ of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules. We review the construction of $X$ from the proof of [19, Theorem 1.10]. Note that $|P|=p$ and $O_{p}(\widetilde{A}_{p})$, the largest $p$-normal subgroup of $\widetilde{A}_{n}$, is the trivial subgroup. The complex $X$ is of the form
@@
where $M$ is in degree zero. Moreover $M$ is the unique (up to isomorphism) summand of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$. Let $Z\to M$ be a projective cover of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $M$. Then $N$ is a direct summand of $Z$ and $\varphi$ is the restriction of the map $Z\to M$ to $N$; moreover, if we write $Z=N\oplus N^{\prime}$, then $N$ and $N^{\prime}$ do not have nonzero isomorphic direct summand.
As in the proof of [21, Lemma 5.2.19], we transport our setup to the group algebra setting. Set
$\mathcal{A}=\langle\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P),(s,s)\rangle\leq\widetilde{S}_{p}\times N_{\widetilde{S}_{p}}(P),$
for some fixed $s\in N_{\widetilde{S}_{p}}(P)\backslash N_{\widetilde{S}_{p}}(P)$. We have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\cong(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}},$
@@
In [21, Proposition 5.2.21], Kleshchev and Livesey proved the following result for Brauer correspondents of blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ with abelian defect groups.
Theorem 6.5 ([21, Proposition 5.2.33]).
$kN_{\widetilde{S}_{n}}(D)e_{\rho,0}$, i.e. $b^{\rho,d}_{k}$, is splendid Rickard equivalent to $kNf$, and $kN_{\widetilde{A}_{n}}(D)e_{\rho,0}$, i.e. $(b^{\rho,d}_{k})_{\bar{0}}$, is splendid Rickard equivalent to $kN_{\bar{0}}f$.
@@
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Then $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e. $b^{\rho,d}_{\mathbb{F}_{p}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N^{\rho,d}f$, and $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e. $(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N_{\bar{0}}^{\rho,d}f$.

### v5

@@ 6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
$G$ $:=\widetilde{S}^{\eta}_{n},$
$L$ $:=\widetilde{S}^{\eta}_{r,p^{d}},$
$N^{\rho,d}$ $:=N_{G}(\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}).$
Note that $N^{\rho,d}$ must permute $V_{1},\cdots,V_{d}$. In particular, $N^{\rho,d}\leq\widetilde{S}^{\eta}_{n}$ is the subgroup generated by $L=\widetilde{S}^{\eta}_{U,V_{1},\cdots,V_{d}}$ and the set $\{T_{w}\mid w\in S_{d}\}$ where the notation $T_{w}$ is from 6.11. Recall the notation in (6.1) that we will also use the subgroups $G_{\bar{0}}=\widetilde{A}_{n}$, $N_{\bar{0}}^{\rho,d}=N^{\rho,d}\cap\widetilde{A}_{n}$, etc. We make the following identifications using (6.3):
$\begin{split}RLe_{z}&\cong R\widetilde{S}^{\eta}_{U}e_{z}\otimes_{R}R\widetilde{S}^{\eta}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{V_{d}}e_{z}.\end{split}$ (6.7)
@@
By Proposition 6.8, we have $f\in R_{p}L\subseteq R_{p}N^{\rho,d}$. Since $N^{\rho,d}$ is generated by $L$ and $\{T_{w}\mid w\in S_{d}\}$, the isomorphism
$RLe_{z}\cong R\widetilde{S}^{\eta}_{r}e_{z}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$
@@
$RN^{\rho,d}e_{z}\cong R\widetilde{S}^{\eta}_{r}e_{z}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{z}\wr_{s}\mathcal{T}^{\eta}_{d,R}),\ T_{i}e_{z}\mapsto(1\otimes\xi_{i}),$
where $T_{i}$’s are chosen to satisfy Lemma 6.12. In particular, we have the isomorphisms of superalgebras
$\begin{split}RLf&\cong R\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{\varnothing,1})^{\otimes d}=B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}\\ RN^{\rho,d}f&\cong R\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}^{\eta}_{d,R})=B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R}).\end{split}$ (6.8)
@@
Since all coefficients involved in the isomorphisms (6.8) are contained in the subring $R_{p}$ of $R$. The isomorphisms (6.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1})^{\otimes d}=B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})=B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}}).\end{split}$
Lemma 6.15 ([23, Lemmas 4.9 and 4.10]).
Let $0<d<p$. As in 6.11, for $i=1,\cdots,d$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $i=1,\cdots,d$ and $w\in S_{d}$, and let
$D:=D_{1}\times\cdots\times D_{d}\leq\widetilde{S}^{\eta}_{V}.$
@@
If $\rho$ is a $d$-Rouquier $\bar{p}$-core, we refer to $B^{\eta,\rho,d}_{R}$ as an RoCK block of weight $d$ over $R$.
For more information on RoCK blocks we refer to [23, §5].
@@
As in Section 6, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$. In this section we continue to adopt the notation in 6.13; we always assume that $d>0$. Let $D$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V}$. Then by Theorem 6.5, $RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RGb=R\widetilde{S}^{\eta}_{n}e_{\rho,d}$, and by [22, Theorem 5.2.10], $RN_{\widetilde{A}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RG_{\bar{0}}b=R\widetilde{A}_{n}e_{\rho,d}$. Note that we interpret $e_{\varnothing,0}$ as $e_{z}$. We set
$\begin{split}b^{\eta,\rho,d}_{R}&:=RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0},~~(b^{\eta,\rho,d}_{R})_{\bar{0}}:=RN_{\widetilde{A}_{n}}(D)e_{\rho,0},\\ b^{\eta,\rho,d}_{R_{p}}&:=R_{p}N_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0},~~(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}N_{\widetilde{A}_{n}}(D)e_{\rho,0}.\end{split}$
Note that $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ (resp. $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $B^{\eta,\rho,d}_{R}$ (resp. $B^{\eta,\rho,d}_{R_{p}}$).
Theorem 7.1 ([22, Proposition 5.2.13]).
Suppose that $0<d<p$. Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $k$-superalgebras
$b^{\eta,\rho,d}_{k}\cong k\widetilde{S}^{\eta}_{U}e_{\rho,0}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}).$ (7.1)
@@
Suppose that $0<d<p$. Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$b^{\eta,\rho,d}_{\mathbb{F}_{p}}\cong\mathbb{F}_{p}\widetilde{S}^{\eta}_{U}e_{\rho,0}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}).$ (7.2)
Proof. By the isomorphism (7.1) and by Proposition 3.4 (i),
$b^{\eta,\rho,d}_{k}\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,\eta}^{k})$ (7.3)
@@
$\begin{split}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}&\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}\\ \mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}&\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}\end{split}$ (7.4)
We realise [22, Lemma 5.2.19] over $\mathbb{F}_{p}$:
@@
Let $P$ be a Sylow $p$-subgroup of $\widetilde{A}_{p}$. There is a splendid Rickard complex of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules that extends to a complex of $(\mathbb{F}_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}^{\eta}_{p}}(P)e_{\varnothing,0})_{C_{2}}$-modules.
Proof. In [20, Theorem 1.10], Kessar and Linckelmann construct a splendid Rickard complex for blocks with a cyclic defect group over a non-splitting field, hence a splendid Rickard complex $X$ of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules. We review the construction of $X$ from the proof of [20, Theorem 1.10]. Note that $|P|=p$ and $O_{p}(\widetilde{A}_{p})$, the largest $p$-normal subgroup of $\widetilde{A}_{n}$, is the trivial subgroup. The complex $X$ is of the form
@@
where $M$ is in degree zero. Moreover $M$ is up to isomorphism the unique nonprojective indecomposable summand of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$. Let $Z\to M$ be a projective cover of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $M$. Then $N$ is a direct summand of $Z$ and $\varphi$ is the restriction of the map $Z\to M$ to $N$; moreover, if we write $Z=N\oplus N^{\prime}$, then $N$ and $N^{\prime}$ do not have nonzero isomorphic direct summand.
As in the proof of [22, Lemma 5.2.19], we transport our setup to the group algebra setting. Set
$\mathcal{A}=\langle\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P),(s,s)\rangle\leq\widetilde{S}^{\eta}_{p}\times N_{\widetilde{S}^{\eta}_{p}}(P),$
for some fixed $s\in N_{\widetilde{S}^{\eta}_{p}}(P)\backslash N_{\widetilde{A}_{p}}(P)$. We have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\cong(\mathbb{F}_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}^{\eta}_{p}}(P)e_{\varnothing,0})_{C_{2}},$
@@
In this setting, we need to show that the complex $X$ from (7.5) extends to a complex of $\mathbb{F}_{p}\mathcal{A}$-modules. Clearly the $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))$-module $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$ extends to an $\mathbb{F}_{p}\mathcal{A}$-module via $(s,s)\cdot x=sxs^{-1}$ for all $x\in e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$. If we write $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}=M\oplus M^{\prime}$, then (by the construction of $X$) $M$ is the nonprojective part and $M^{\prime}$ is the projective part of $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$. Hence both $M$ and $M^{\prime}$ are $C_{2}$-invariant; that is, $(s,s)M\cong M$ and $(s,s)M^{\prime}\cong M^{\prime}$ as $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))$-modules. By [8, Theorem 3.11], $M$ and $M^{\prime}$ extend to $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$-modules. Let $\tilde{M}$ be such an extension of $M$ to $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$. Let $\tilde{Z}$ be a projective cover of the $\mathbb{F}_{p}\mathcal{A}$-module $\tilde{M}$. Since $2$ is invertible in $\mathbb{F}_{p}$, we have
$\tilde{Z}/J(\mathbb{F}_{p}\mathcal{A})\tilde{Z}\cong\tilde{M}/J(\mathbb{F}_{p}\mathcal{A})\tilde{M}\cong\tilde{M}/J(\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P)))\tilde{M}\cong M/{\rm rad}(M)$
@@
In [22, Proposition 5.2.21], Kleshchev and Livesey proved the following result for Brauer correspondents of blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ with abelian defect groups.
Theorem 7.5 ([22, Proposition 5.2.33]).
Suppose that $0<d<p$. Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$. Then $kN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$, i.e. $b^{\eta,\rho,d}_{k}$, is splendid Rickard equivalent to $kNf$, and $kN_{\widetilde{A}_{n}}(D)e_{\rho,0}$, i.e. $(b^{\eta,\rho,d}_{k})_{\bar{0}}$, is splendid Rickard equivalent to $kN_{\bar{0}}f$.
@@
Suppose that $0<d<p$. Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$. Then $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e. $b^{\eta,\rho,d}_{\mathbb{F}_{p}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N^{\rho,d}f$, and $\mathbb{F}_{p}N_{G_{\bar{0}}}(D)e_{\rho,0}$, i.e. $(b^{\eta,\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N_{\bar{0}}^{\rho,d}f$.